OpenStudy (noahbred):

FAN and MEDAL Which expression shows how 8 • 54 can be rewritten using the distributive property? A. 4 • 50 + 4 • 4 B. 8 • 50 + 8 • 4 C. 8 • 50 + 8 • 54 D. 50 • 8 + 50 • 4

OpenStudy (texaschic101):

8 * 50 + 8 * 4

OpenStudy (photon336):

interesting so essentially it seems like it was written like this. \[58 = (50+8) \] \[8(50+8) = 8*50+8*8 \]

OpenStudy (anonymous):

8*54=8(4+50)=?